# Crepant Transformation correspondence For Toric Stack Bundles

### v1

@@ 3.3. Fourier-Mukai Transformation
Since $D_{j_{+}}\cdot e$ is a positive integer, this factor is a polynomial in $t^{-1}$ and its restriction to $\delta$ vanishes. So the only possibility for non-zero restriction is $\delta_{+}$ with $\delta_{+}|\delta_{-}$. Apply localization formula one more time, we obtain
$\mathbb{FM}(e_{\delta_{-},\varrho})=\frac{1}{l}\sum_{t\in\mathcal{T}}\frac{1-S^{+}_{j_{-}}}{1-t^{-1}}L_{+}(\hat{\varrho})t^{\hat{\varrho}\cdot e}\prod_{i\notin\delta_{-}}(1-t^{-k_{i}}S^{+}_{i}).$
@@ 3.4. Comparison
First we check commutativity on basis $K^{0}_{T}(B)$. Note that $\tilde{\text{ch}}$ maps $K^{0}_{T}(B)$ into $H^{\bullet}(B)\otimes R_{T}$ and everything is fixed under $\mathbb{FM}$ and $\mathbb{U}_{H}$ in this case.
@@
For $\delta_{-}\notin\mathcal{A}_{-}\cap\mathcal{A}_{+}$, we calculate them explicitly. First note that we have
$\tilde{\text{ch}}(L_{\pm}(\hat{\varrho}))$ $=\bigoplus_{f\in\mathbb{K}_{\pm}/\mathbb{L}}e^{2\pi i\hat{\varrho}\cdot f}e^{\theta_{\pm}(\hat{\varrho})}\mathbf{1}_{f}.$
@@
$\tilde{\text{ch}}(S^{\pm}_{j})$ $=\bigoplus_{f\in\mathbb{K}_{\pm}/\mathbb{L}}e^{-2\pi iD_{j}\cdot f}e^{-U_{j}}\mathbf{1}_{f}$
@@
$\mathbb{FM}(e_{\delta_{-},\varrho})=\sum_{t\in\mathcal{T}}\left(\frac{1-S^{+}_{j_{-}}}{l(1-t^{-1})}\prod_{i\notin\delta_{-},D_{i}\cdot e<0}\frac{1-S^{+}_{i}}{1-t^{-D_{i}\cdot e}S^{+}_{i}}\right)\left(L_{+}(\hat{\varrho})t^{\hat{\varrho}\cdot e}\prod_{i\notin\delta_{-}}(1-t^{-D_{i}\cdot e}S^{+}_{i})\right).$
@@
$i^{\star}_{(\delta_{+},f_{+})}\tilde{\text{ch}}(\mathbb{FM}(e_{\delta_{-},\varrho}))=\sum_{(\delta^{\prime}_{-},f_{-})|(\delta_{+},f_{+})}C^{\delta^{\prime}_{-},f_{-}}_{\delta_{+},f_{+}}(i^{\star}_{(\delta^{\prime}_{-},f_{-})}\tilde{\text{ch}}(e_{\delta_{-},\varrho}))$
First using Lemma 3.2 we have
@@
$=i^{\star}_{(\delta_{-},f_{-})}\tilde{\text{ch}}(L_{-}(\hat{\varrho})),$
@@
$i^{\star}_{(\delta_{+},f_{+})}\tilde{\text{ch}}\left(\frac{1-S^{+}_{j_{-}}}{l(1-t^{-1})}\prod_{i\notin\delta_{-},D_{i}\cdot e<0}\frac{1-S^{+}_{i}}{1-t^{-D_{i}\cdot e}S^{+}_{i}}\right)$
@@
×$\times e^{-U_{j_{-}}(\delta_{+})/2-\pi iD_{j_{-}}\cdot f_{+}+\frac{1}{2l}(U_{j_{-}}(\delta_{+})+2\pi iD_{j_{-}}(f_{+}-f_{-}))+\sum_{i\notin\delta_{-},D_{i}\cdot e<0}(-U_{i}(\delta_{+})/2-\pi iD_{i}\cdot f_{+}+U_{i}(\delta_{-})/2+\pi iD_{i}\cdot f_{-})}$
@@
$i^{\star}_{(\delta_{+},f_{+})}\tilde{\text{ch}}(\mathbb{FM}(e_{\delta_{-},\varrho}))$ $=\sum_{t\in\mathcal{T}}i^{\star}_{(\delta_{+},f_{+})}\tilde{\text{ch}}\left(\frac{1-S^{+}_{j_{-}}}{l(1-t^{-1})}\prod_{i\notin\delta_{-},D_{i}\cdot e<0}\frac{1-S^{+}_{i}}{1-t^{-D_{i}\cdot e}S^{+}_{i}}\right)$
×$\times i^{\star}_{(\delta_{+},f_{+})}\tilde{\text{ch}}\left(L_{+}(\hat{\varrho})t^{\hat{\varrho}\cdot e}\prod_{i\notin\delta_{-}}(1-t^{-D_{i}\cdot e}S^{+}_{i})\right)$

### v2

@@ 3.3. Fourier-Mukai Transformation
Since $D_{j_{+}}\cdot e$ is a positive integer, this factor is a polynomial in $t^{-1}$ and its restriction to $\delta$ vanishes. So the only possibility for non-zero restriction occurs for $\delta_{+}$ with $\delta_{+}|\delta_{-}$. Applying localization formula one more time, we obtain
[…]
@@ 3.4. Comparison
First, we check commutativity on basis $K^{0}_{T}(B)$. Note that $\tilde{\text{ch}}$ maps $K^{0}_{T}(B)$ to $H^{\bullet}(B)\otimes R_{T}$ and everything is fixed under $\mathbb{FM}$ and $\mathbb{U}_{H}$ in this case.
@@
For $\delta_{-}\notin\mathcal{A}_{-}\cap\mathcal{A}_{+}$, we calculate them explicitly. First, note that we have
$)\displaystyle\tilde{\text{ch}}(L_{\pm}(\hat{\varrho}$)) $=\bigoplus_{f\in\mathbb{K}_{\pm}/\mathbb{L}}e^{2\pi i\hat{\varrho}\cdot f}e^{\theta_{\pm}(\hat{\varrho})}\mathbf{1}_{f}.$
@@
$\tilde{\text{ch}}(S^{\pm}_{j})$ $=\bigoplus_{f\in\mathbb{K}_{\pm}/\mathbb{L}}e^{-2\pi iD_{j}\cdot f}e^{-U_{j}}\mathbf{1}_{f},$
@@
[…]
@@
$i^{\star}_{(\delta_{+},f_{+})}\tilde{\text{ch}}(\mathbb{FM}(e_{\delta_{-},\varrho}))=\sum_{(\delta^{\prime}_{-},f_{-})|(\delta_{+},f_{+})}C^{\delta^{\prime}_{-},f_{-}}_{\delta_{+},f_{+}}(i^{\star}_{(\delta^{\prime}_{-},f_{-})}\tilde{\text{ch}}(e_{\delta_{-},\varrho})).$
First, using Lemma 3.2 we have
@@
$),\displaystyle=i^{\star}_{(\delta_{-},f_{-})}\tilde{\text{ch}}(L_{-}(\hat{\varrho}$)),
@@
[…]
@@
$\times e^{-U_{j_{-}}(\delta_{+})/2-\pi iD_{j_{-}}\cdot f_{+}+\frac{1}{2l}(U_{j_{-}}(\delta_{+})+2\pi iD_{j_{-}}(f_{+}-f_{-}))+\sum_{i\notin\delta_{-},D_{i}\cdot e<0}(-U_{i}(\delta_{+})/2-\pi iD_{i}\cdot f_{+}+U_{i}(\delta_{-})/2+\pi iD_{i}\cdot f_{-})}$
@@
[…]